OpenStudy (anonymous):

Please Help! Need One Last Question! A tugboat goes 160 miles upstream in 20 hours. The return trip downstream takes 10 hours. Find the speed of the tug boat without a current and the speed of the current.

OpenStudy (aravindg):

so tell me what is speed during upstream journey

OpenStudy (anonymous):

160miles? @AravindG

OpenStudy (aravindg):

speed=distance/time

OpenStudy (anonymous):

??????? still confused can you write the equation out please..... @AravindG

OpenStudy (aravindg):

what is distance=160 miles time =20 hrs speed =?

OpenStudy (anonymous):

would it be 8 @AravindG

OpenStudy (aravindg):

yep :)

OpenStudy (anonymous):

Is that the answer to the speedof the tugboat with out a current? @AravindG

OpenStudy (aravindg):

no can u find speed of return trip now?

OpenStudy (anonymous):

Would the answer be 16 for the current? @AravindG

OpenStudy (aravindg):

hey i think we should do it by another method

OpenStudy (anonymous):

so those answers are incorrect? @AravindG

OpenStudy (aravindg):

Let s= the speed of the tugboat without a current Let c= the speed of the current Going upstream: 160=(s-c) *20 Going downstream: 160=(s+c)*10

OpenStudy (aravindg):

now find s+c and s-c from these 2 eqns

OpenStudy (anonymous):

For the first equation i multiply 20 to both s and c? @AravindG

OpenStudy (aravindg):

for first eqn divide by 20

OpenStudy (anonymous):

so is it 8=(s-c) @AravindG

OpenStudy (aravindg):

ya what abt second eqn ?

OpenStudy (anonymous):

16=(s+c) @AravindG

OpenStudy (aravindg):

yep :)

OpenStudy (aravindg):

now add both these to get c

OpenStudy (aravindg):

then u will get speed of current !

OpenStudy (anonymous):

so 24? @AravindG

OpenStudy (aravindg):

try again ?

OpenStudy (anonymous):

2? @AravindG

OpenStudy (aravindg):

try again ?

OpenStudy (anonymous):

Ya i dont know @AravindG

OpenStudy (aravindg):

(s-c) = 8 + (s+c) =16 __________

OpenStudy (aravindg):

sorry u get s vakue when u add not c ...anywas find s

OpenStudy (aravindg):

so can u add it now ?

OpenStudy (anonymous):

8? @AravindG or 24

OpenStudy (aravindg):

can u show ur addition using draw button below ?

OpenStudy (anonymous):

i mean 12 @AravindG

OpenStudy (aravindg):

\[\checkmark \]

OpenStudy (aravindg):

now substitute s value in any of the above eqns to get c :)

OpenStudy (anonymous):

4? @AravindG

OpenStudy (aravindg):

\[\checkmark \]

OpenStudy (anonymous):

THANK YOU VERY MUCH! i have one more question can you help me on my next problem?

OpenStudy (anonymous):

@AravindG

OpenStudy (aravindg):

dont u need speed of tug boat without current ?

OpenStudy (anonymous):

yes i do.... how do we figure that out?? @AravindG

OpenStudy (aravindg):

my mistake :P speed of tug boat in still water is s=12 :P

OpenStudy (aravindg):

lols

OpenStudy (anonymous):

ok aha thank you again @AravindG

OpenStudy (aravindg):

yw :)